### v2

# Electron-phonon couplings in polymorphous crystals

Abstract
Positional polymorphism in solids refers to locally disordered unit cells that, on average, reproduce the high-symmetry structures observed in diffraction experiments. Standard theories of electron-phonon interactions fail to describe the temperature-dependent electronic structure of such polymorphous systems. Hybrid halide perovskites are a prime example, where configurational entropy from both polymorphism and molecular disorder plays a central role. Here we generalize the special displacement method to polymorphous crystals, providing an efficient ab initio framework for electron-phonon couplings without resorting to molecular dynamics. We resolve long-standing discrepancies in hybrid halide perovskite physics, including temperature-dependent anharmonic phonons and band gaps. Our approach provides a practical route to link local disorder, configurational entropy, and electron-phonon interactions, with applicability across diverse material classes, from optoelectronics and ferroelectrics to thermoelectrics.
Recently, the concept of correlated local disorder (positional polymorphism) has gained significant attention across a broad range of materials Zacharias and Even (2025); Zhao et al. (2020); Wang et al. (2020); Fabini et al. (2020); Wang et al. (2021); Goesten et al. (2022); Zacharias et al. (2023a); Dirin et al. (2023); Sabisch et al. (2025); Wang et al. (2025). In perovskites, local disorder has been observed experimentally since the 1960s Comes et al. (1968), and later confirmed through x-ray and neutron scattering as well as nuclear magnetic resonance measurements Itoh et al. (1994); Mashiyama et al. (1998); Kiat et al. (2000); Worhatch et al. (2008); Beecher et al. (2016); Senn et al. (2016); Laurita et al. (2017); Sutton et al. (2018); Ferreira et al. (2020); Doherty et al. (2021); Dirin et al. (2023); Morana et al. (2023); Weadock et al. (2023); Yazdani et al. (2023); He et al. (2024); Sabisch et al. (2025); Dubajic et al. (2025). A prominent class of such materials is hybrid halide perovskites Mashiyama et al. (1998); Zhao et al. (2020), which exhibit high efficiency and favorable optoelectronic properties, making them excellent candidates for energy applications such as photovoltaics and light-emitting devices Kojima et al. (2009); Snaith (2013); Lin et al. (2018); Sidhik et al. (2022); Metcalf et al. (2023); Correa-Baena et al. (2017); Mao et al. (2018). To realize the full potential of halide perovskites, a fundamental understanding of positional polymorphism, configurational entropy, and their interplay with the electronic, vibrational, and electron-phonon properties is essential.
Hybrid ABX3 halide perovskites are composed of an organic molecule (A) and an inorganic network formed by metal cations (B) octahedrally coordinated by halide anions (X) [Figs. 1(a,b)]. One central assumption in static density functional theory (DFT) calculations for tetragonal or cubic hybrid halide perovskites is the use of a periodic high-symmetry (monomorphous) metal-halide network Even et al. (2013a, b); Amat et al. (2014); Saidi et al. (2016); Whalley et al. (2017); Mao et al. (2018); Traoré et al. (2019); Poncé et al. (2019); Ghaithan et al. (2020). The orientation of the molecule A is typically randomly chosen or set to towards an unrealistic static direction. Critical issues related to the use of a high-symmetry inorganic network include (i) neglecting important corrections to the electronic structure Zhao et al. (2020), (ii) the breakdown of the standard harmonic approximation to describe lattice dynamics Zacharias et al. (2023b), (iii) neglecting energetically more stable configurations and thus overlooking important structural features impacting chemical bonding and optoelectronic properties Katan et al. (2018); Quarti et al. (2024); Garba et al. (2025), and (iv) ignoring important modifications to electron-phonon coupling (EPC) Zacharias et al. (2023a).
Figure 1: (a-f) Ball-stick models of cubic FAPbI3 and MASnI3 representing monomorphous (a,b), reference (c,d), and polymorphous (e,f) structures. Structures in (a,b) refer to unit cells (12 atoms), in (c,d,e,f) to $\times2$$\times$2 supercells. (g,h) Calculated temperature-dependent band gaps (coloured discs) of Pb (g) and Sn (h) cubic hybrid halide perovskites including electron-phonon coupling (EPC) and thermal expansion (TE) effects; EPC is incorporated via the ASDM using ZG polymorphous structures and TE contribution is evaluated by allowing lattice expansion according to the experimental expansion coefficients Zacharias and Even (2026). Experimental gaps (triangles) are from Refs. [Milot et al., 2015; Kontos et al., 2018; Parrott et al., 2016; Mannino et al., 2020; Chen et al., 2020; Kahmann et al., 2020]. ASDM calculations refer to $\times4$$\times$4 supercells (768 atoms) and the DFT-PBEsol including SOC. Data are shifted close to PBE0 corrections (see Table I of Ref. [Zacharias and Even, 2026]) to match experiment. PBE0 corrections to electron-phonon renormalized band gaps of halide perovskites are negligible Zacharias et al. (2023a).
An approach for more accurate static DFT calculations using locally disordered (polymorphous) perovskites has been proposed Zhao et al. (2020); Wang et al. (2021). It has been shown that accounting for symmetry-breaking domains in tetragonal and cubic perovskites yields considerably improved agreement with measurements on pair distribution functions, band gaps, and effective masses Zhao et al. (2020); Wang et al. (2021); Zacharias et al. (2023a). Although molecular dynamics (MD) Quarti et al. (2015, 2016); Even et al. (2016); Wiktor et al. (2017); Carignano et al. (2017); Lahnsteiner and Bokdam (2022); Diez-Cabanes et al. (2023); Weadock et al. (2023); Seidl et al. (2023); He et al. (2024); Dai et al. (2024a) can capture local disorder through long equilibration times, a direct distinction between local disorder and thermal vibrations is inevitably missed.
Recently, we have introduced a bottom-up first-principles approach Zacharias et al. (2023a), showing that local disorder in inorganic halide perovskites (A=Cs) yields dynamically stable and strongly interacting phonons, and affects profoundly anharmonic EPCs. This polymorphous framework can be seen as a quasi-static approximation of the slow relaxation dynamics of the structure. In this Letter, we explore the more challenging and technologically relevant cases of hybrid halide perovskites and compare with their inorganic counterparts. Together with Ref. [Zacharias and Even, 2026], we propose an efficient approach to evaluate phonons and temperature-dependent band gaps of polymorphous cubic APbBr3, APbI3, and ASnI3, where A = methylammonium (MA) or formamidinium (FA). Our high-throughput calculations agree with experiments for all cases, as shown in Fig. 1, opening the way for effective predictions of anharmonic electron-phonon properties of polymorphous systems with configurational entropy.
DFT calculations were performed using Quantum Espresso Giannozzi et al. (2009, 2017) in conjunction with optimized norm-conserving Vanderbilt pseudopotentials Hamann (2013); van Setten et al. (2018) and the Perdew-Burke-Ernzerhof exchange-correlation functional revised for solids (PBEsol) Perdew et al. (2008); full computational details are available in Ref. [Zacharias and Even, 2026]. In Figs. 1(a)-(f), we show the geometries of the cubic FAPbI3 and MASnI3. We consider three cases: (i) the monomorphous structure represented by a minimal unit cell [i.e. the fomula unit (f.u.)] where the atoms of the inorganic network reside at their high-symmetry positions and the FA or MA molecules point towards a single orientation [Figs. 1(a) and (b)], (ii) the reference structure represented by a supercell where the atoms of the inorganic network are fixed at their high-symmetry positions and the FA or MA molecules are relaxed starting from random orientations [Figs. 1(c) and (d)], (iii) the polymorphous structure represented by a supercell where the atoms of the inorganic network and FA or MA molecules are relaxed starting from the reference structure [Figs. 1(e) and (f)]. During relaxations, the lattice constants are fixed at their experimental values Zacharias and Even (2026). We consider reference configurations since changes in the electronic structure and phonons are sensitive to the net dipole moments induced by single cation orientations and to ensure a consistent comparison with polymorphous structures. In Cs-based compounds, the reference is identical to the monomorphous structure so no special treatment is required. Additionally, to assess the electronic structure and thermal effects variability of the polymorphous networks originating from configurational entropy Quarti et al. (2015); Motta et al. (2015), we generate 10 distinct polymorphous structures and take Boltzmann-weighted averages relying on the system’s free energy Zacharias and Even (2026). Convergence tests of the band gap as a function of the number of polymorphous configurations are provided in Ref. [Zacharias and Even, 2026].
To compute thermal band gap renormalization, we consider both EPC and lattice thermal expansion (TE) so that the band gap temperature coefficient is given by Villegas et al. (2016):
${dE_{\rm g}}/{dT}={dE_{\rm g}}/{dT}|_{\rm EPC}+{dE_{\rm g}}/{dT}|_{\rm TE}.$ (1)
We use temperature-dependent phonons computed within the self-consistent phonon theory as implemented in the anharmonic special displacement method Zacharias et al. (2023b) (ASDM). This requires computing interatomic force constants iteratively of the corresponding thermally displaced reference structures until convergence in the phonons and system’s free energy Zacharias and Even (2026). The computed dynamically stable phonons allow us to include phonon anharmonicity in nonperturbative calculations of EPC via the special displacement method Zacharias and Giustino (2020); Lee et al. (2023); Zacharias et al. (2023a). That is we apply special Zacharias-Giustino (ZG) displacements Zacharias and Even (2026) starting from the polymorphous structures. To account for TE, we expand the lattice based on the experimental expansion coefficients Kawamura et al. (2002); Mashiyama et al. (2003); Stoumpos et al. (2013); Rakita et al. (2015); Whitfield et al. (2016); Dang et al. (2016); Fabini et al. (2016); Schueller et al. (2017); Kontos et al. (2018); Marronnier et al. (2018); Keshavarz et al. (2019); Handa et al. (2020); He et al. (2021), as decribed in Ref. [Zacharias and Even, 2026].
In Figs. 1(g,h), we compare our calculations for temperature-dependent band gaps of polymorphous cubic MAPbBr3, MAPbI3, FAPbBr3, FAPbI3, FASnI3, and MASnI3 with experiments Milot et al. (2015); Kontos et al. (2018); Parrott et al. (2016); Mannino et al. (2020); Chen et al. (2020); Kahmann et al. (2020). We include data for CsPbBr3 and CsSnI3 for completeness Zacharias and Even (2026). Calculations and experiments agree for all compounds, showcasing the broad applicability of our approach in describing anharmonic electron-phonon renormalized band structures. We note that an empirical shift is applied only to align the absolute band gap values with experiment. Quantitative agreement of temperature-dependent trends is evidenced by our calculated $dE_{\rm g}/dT$ reported in Table 1, in which we further provide data for CsPbI3 Zacharias and Even (2026). We also show the EPC contribution to the gap renormalization for all materials, varying between 27 – 97%. For the majority of compounds, EPC plays a dominant role in band gap renormalization, challenging the prevailing view that TE and EPC contribute equally in lead halide perovskites Francisco-López et al. (2019); Rubino et al. (2020). A clear trend is observed in the variation of the EPC contribution across all compounds, with its percentage increasing as the A-site cation changes from FA to MA and then to Cs. For example, in FAPbBr3 and FASnI3, EPC contributes only 27% and 34%, respectively. This lower contribution is not surprising, as their volumetric TE coefficients are among the highest reported for halide perovskites Schueller et al. (2017). Sn-based compounds exhibit a larger $dE_{\rm g}/dT$, which we attribute to positional polymorphism enhancing band gap renormalization via TE. This effect is driven by the stronger lone pair activity Balvanz et al. (2024) of Sn that promotes Sn-I-Sn bending under TE Zacharias and Even (2026). We note that small deviations from experiment arise from approximations in our framework (e.g. constant-temperature phonons or neglect of non-adiabatic and many-body effects Antonius et al. (2014); Miglio et al. (2020)), and from limited experimental data in some cases such as CsPbI3.
Table 1: Band gap temperature coefficient (${dE_{\rm g}}/{dT}$) of cubic halide perovskites. r- and p- stand for reference and polymorphous. Experimental values are extracted by linear fits Zacharias and Even (2026).
${dE_{\rm g}}/{dT}\big|_{\rm expt.}$   ${dE_{\rm g}}/{dT}$ ${dE_{\rm g}}/{dT}\big|_{\rm EPC}$    EPC
($10^{-4}$ eV/K)    Pct.
r-MAPbBr3 2.8 5.8 4.0 71%
p-FAPbBr3 4.9 3.6 1.2 34%
p-MAPbBr3 2.8 2.9 2.0 70%
p-CsPbBr3 3.4 2.6 2.4 94%
p-FAPbI3 4.3 4.6 3.2 70%
p-MAPbI3 2.6 3.1 2.7 88%
p-CsPbI3 8.4 1.2 1.2 97%
p-FASnI3 10.6 8.0 2.2 27%
p-MASnI3 4.5 5.7 3.5 61%
p-CsSnI3 4.7 3.9 3.3 84%
Now we analyze the effect of local disorder in the electronic structure of hybrid halide perovskites without thermal effects. In Figs. 2(a) and (b), we compare the electron spectral functions (color maps) obtained for the polymorphous cubic FAPbI3 and MASnI3 with the band structures (black lines) calculated for their monomorphous networks. Spectral functions were calculated using band unfolding Popescu and Zunger (2012) including spin orbit coupling (SOC) as implemented in bands_unfold.x of the EPW/ZG module Zacharias and Giustino (2020); Lee et al. (2023). Local disorder leads to significant changes in the electronic structure, including enhanced band broadenings and gap openings. Particularly, our calculations for the polymorphous cubic FAPbI3 yield an average hole and electron effective mass enhancements Giustino (2017); Zacharias and Even (2026) of $\lambda_{\rm h}=2.3$ and $\lambda_{\rm e}=2.0$, and a gap opening of $\Delta E_{\rm g}=0.24$ eV compared to the reference structure. The latter value compares well with $\Delta E_{\rm g}=0.29$ eV with respect to the monomorphous structure, reported in Ref. [Zhao et al., 2020]. The corresponding results for polymorphous MASnI3 are $\Delta E_{\rm g}=0.37$ eV, $\lambda_{\rm h}=1.2$, and $\lambda_{\rm e}=3.0$.
Figure 2: (a,b) Electron spectral functions (color maps) of the polymorphous cubic FAPbI3 (a) and MASnI3 (b). Black dispersions represent the band structure of the monomorphous networks. Calculations are at the PBEsol+SOC level and thus underestimate the band gap. Combining hybrid functionals with SOC and polymorphous structures reproduce experiments well Zacharias and Even (2026); Garba et al. (2025). (c,d) Average band gap increase due to positional polymorphism $\Delta E_{g}$ (c) and tolerance factor $t$ (d) versus the average metal-halide-metal angle decrease $\Delta\theta_{\rm B-X-B}$ calculated for 10 distinct polymorphous configurations. Solid lines are fits obtained by linear regression and shaded regions represent three standard deviations on either side of the lines. The error bars represent the standard deviation across 10 configurations used for each material.
Figure 2(c) show the dependence of the band gap increase due to positional polymorphism $\Delta E_{\rm g}$ in hybrid halide perovskites on the average bond angle decrease $\Delta\theta_{\rm B-X-B}$. For completeness, we also include data for Cs- and Cl-based halide perovskites from Ref. [Zacharias and Even, 2026]. The measure Baur (1974); Filip et al. (2014) $\Delta\theta_{\rm B-X-B}$ is strongly correlated with bond length changes Zacharias and Even (2026) and thereby is chosen as the main descriptor for quantifying local disorder with respect to the high-symmetry inorganic network. Our calculations and analysis yield $\Delta E_{\rm g}$ in the range 198–616 meV and a slope of 22 meV/(), revealing a strong correlation between $\Delta E_{\rm g}$ and bond angles change with a Pearson correlation of $r=0.96$. This is expected, as deviations in bond angles from their ideal values lead to a reduction in the antibonding coupling between metal and halogen orbitals Knutson et al. (2005); Filip et al. (2014); Dar et al. (2016); Meloni et al. (2016); Zacharias and Even (2026). The high $R^{2}$ of 0.92 show that this distortion measure can effectively explain the gap opening due to positional polymorphism. Furthermore, the band gap variability, computed for 10 configurations and shown as error bars [Fig. 2(c)], is larger for MA-based compounds due to the larger dipole moments of the MA compared to the FA cations Carignano et al. (2017); Maheshwari et al. (2019). Cl compounds, e.g. MAPbCl3, exhibit the largest band gap variability under local disorder because the smaller, more electronegative Cl ions create stronger Pb-Cl bonds, making the band structure highly sensitive to structural distortions.
Compared to their reference structures, polymorphous cubic FA, MA, and Cs-based compounds exhibit total energy lowerings of 50–73 meV/f.u., 53–93 meV/f.u., and 61–121 meV/f.u., respectively Zacharias and Even (2026). This trend highlights the role of A-cation size in stabilizing locally disordered phases, with smaller cations facilitating structural distortions Kieslich et al. (2014); Li et al. (2022). This is further illustrated in Figs. 2(c) and (d), where bond angle variations depend strongly on the cations and the tolerance factors. We evaluate the tolerance factors ($t$) using ionic radii of 1.19, 1.15, 2.2, 1.96, 1.81, 1.88, 2.2, and 2.5 Å  for Pb+2, Sn+2, I-, Br-, Cl-, Cs+, MA+, and FA+ Shannon (1976); Kieslich et al. (2015), respectively. Our analysis suggests that smaller tolerance factors allow for greater local disorder, and thus larger band gap openings. For instance, the relatively low tolerance factors of Cs-based halide perovskites (t=0.85t=0.850.870.87) facilitate octahedral distortion due to the absence of larger organic A-site cations Kieslich et al. (2014); Filip and Giustino (2018); Li et al. (2022). This structural flexibility promotes a higher local disorder, which leads to significant band gap renormalizations exceeding 0.55 eV. Similarly, MA-based compounds exhibit a higher degree of polymorphism, smaller $t$, and larger $\Delta E_{\rm g}$ than FA-based compounds, attributable to the smaller size of MA.
Figure 3: (a,b,c) Temperature-dependent band gap of cubic MAPbBr3 calculated considering (a) electron-phonon coupling (EPC), (b) lattice thermal expansion (TE), and (c) both effects combined. Calculated data are shifted by 1.38 eV (orthorhombic), 1.79 eV (monomorphous cubic) and 1.44 eV (polymorphous cubic) to match the experiment Mannino et al. (2020) (grey triangles) at $T=275$ K. (d) Eliashberg function showing the one-phonon energy-resolved band gap renormalization of MAPbBr3 at $T=300$ K. (e) Fan-Migdal (FM) and Debye-Waller (DW) self-energy corrections to the VBM of the reference and polymorphous cubic FAPbI3 due to a bending mode. (f,g) Linear variations in the charge density at the VBM obtained for a single motif of the reference (f) and polymorphous (g) networks. Arrows representing the bending mode are scaled by 50. The isosurface value $u$ is indicated.
At this point, we emphasize that local disorder in hybrid halide perovskites appears to be a key factor in determining the band gap order between FA- and MA-based compounds. For example, as shown in Fig. 1(g), MA-based cubic lead halide perovskites have a larger band gap compared to their FA-based counterparts within the same temperature range. The difference can be attributed to the band gap widenings reported in Fig. 2(c). In the case of tin halide perovskites, the situation is reversed, a trend that agrees with experiments, as shown in Fig. 1(h). This observation is explained, e.g., by the larger TE coefficient of FASnI3 Schueller et al. (2017), together with the greater propensity for lone pair expression in sterically constrained FA-based compounds, which enhance thermal-induced band gap renormalization.
We next elucidate the mechanisms determining the band gap temperature coefficient of halide perovskites. Figure 3(a) shows the electron-phonon renormalized band gap of MAPbBr3 as a function of temperature calculated for the ZG reference (red) and ZG polymorphous (blue) cubic structures, both generated using stable anharmonic phonons. Local disorder significantly affects EPC, leading to a slower band gap variation with temperature. On the same plot, we include calculations for the orthorhombic structure (orange) using harmonic phonons. Attempting to mimic positional polymorphism at high temperatures using the orthorhombic phase yields improved results but does not fully accurately capture the effect of EPC. Notably, polymorphism also affects the band gap widening due to TE, as shown in Fig. 3(b); a point that has been overlooked in previous calculations Foley et al. (2015); Saidi et al. (2016); Francisco-López et al. (2019); Ning et al. (2022). Accounting for local disorder in our calculations reduces the band gap opening due to TE by 50% and the slope $dE_{\rm g}/dT|_{\rm TE}$ alone underestimates significantly the experimental data Mannino et al. (2020).
In Fig. 3(c), we show that combining EPC and lattice expansion in our calculations for the polymorphous cubic MAPbBr3, the total $dE_{\rm g}/dT$ yields excellent agreement with experiment. In contrast, $dE_{\rm g}/dT$ is overestimated by a factor of two when the ZG reference network is used. A quantitative comparison of $dE_{\rm g}/dT$ alongside the percentage contribution of EPC for reference and polymorphous MAPbBr3 is presented in Table 1.
Figure 3(d) shows the one-phonon energy-resolved contribution to the band gap renormalization of MAPbBr3 at $T=300$ K calculated using the Eliashberg function: $g^{2}F_{n}(\omega,T)=\sum_{\nu}\Delta\varepsilon_{n,\nu}(T)\delta(\hbar\omega-\hbar\omega_{\nu}).$ Here, $\hbar\omega_{\nu}$ is the phonon energy of mode $\nu$, $\Delta\varepsilon_{n,\nu}(T)=\frac{1}{2}\frac{\partial^{2}\varepsilon_{n}}{\partial x_{\nu}^{2}}\sigma^{2}_{\nu,T}$ is the mode-resolved energy renormalization of a state $\psi_{n}$ up to second order, and $\sigma^{2}_{\nu,T}$ is the associated mean-square displacement Zacharias and Giustino (2016). $\Delta\varepsilon_{n,\nu}(T)$ is calculated by finite differences using single mode displaced configurations generated by ZG.x Lee et al. (2023). Our results for the reference and polymorphous structures reveal that low energy optical vibrations in the window 2–10 meV and 4–10 meV, respectively, dominate the phonon-induced band gap widening, in agreement with interpretations of previous works Zacharias et al. (2023a); Yazdani et al. (2023); Zhu et al. (2024); Biswas et al. (2024). The modes associated with this energy window have a bending or rocking character associated with octahedral tilting Pérez-Osorio et al. (2015). Higher frequency modes are mostly associated to B-X stretching or molecular librations. Positional polymorphism strongly suppress the positive EPC contribution in the window 2–10 meV, while enhancing the negative contributions in the window 10–20 meV, resulting in significant reduction in the overall gap opening.
To elucidate the origin of this reduction, we investigate the band gap renormalization due to a mode with bending character. If we consider the Kohn-Sham Hamiltonian $H_{\rm KS}$, express $\varepsilon_{n}=\braket{\psi_{n}|H_{\rm KS}|\psi_{n}}$, take the second derivative with $x_{\nu}$, and apply the Hellmann-Feynman theorem twice Poncé et al. (2014) we obtain:
$\hskip-11.38092pt\frac{\partial^{2}\varepsilon_{n}}{\partial x_{\nu}^{2}}=\int d{\bf r}\frac{\partial V_{\rm KS}}{\partial x_{\nu}}\frac{\partial|\psi_{n}({\bf r})|^{2}}{\partial x_{\nu}}+\int d{\bf r}\frac{\partial^{2}V_{\rm KS}}{\partial x_{\nu}^{2}}|\psi_{n}({\bf r})|^{2}.$ (2)
Here, $V_{\rm KS}$ is the Kohn-Sham potential and the integration is taken over all electron coordinates $\bf r$}. To arrive at this expression we also neglected the non-local contribution to $V_{\rm KS}$ Starace (1971). The first and second terms on the right hand side of Eq. (2) are the Fan-Migdal (FM) and Debye-Waller (DW) EPC coefficients Giustino (2017).
Figure 3(e) shows the FM and DW self-energy corrections to the valence band maximum (VBM) induced by nuclei vibrations of amplitude 0.02 Å along a bending mode. We present results for the reference and polymorphous networks of FAPbI3. In both cases, the negative FM dominates over the positive DW correction leading to an overall energy decrease of the VBM of 8.2 meV and 1.5 meV for the reference and polymorphous networks, respectively. This contributes to the band gap opening in both cases. The relatively large difference in the VBM energy decrease is associated to the strong reduction of the FM term (62%) when we employ a polymorphous network. This result can be explained by inspecting the changes of the electronic charge distribution at the VBM in response to the movement of the nuclei along the bending mode (i.e. $\partial|\psi({\bf r})|^{2}/\partial x_{\nu}$), shown in Figs. 3(f) and (g). The charge density in the reference structure exhibits a higher response to the nuclear vibrations than that in the polymorphous structure, leading essentially to a much larger self-energy correction. The reduced response in the polymorphous structure arises from the deviation of Pb-I-Pb bond angles from 180, which disrupts orbital overlap and localizes the electronic states, thereby weakening the electron-phonon coupling. On the other hand, the DW corrections to the VBM of the reference and polymorphous structures are nearly the same [Fig. 3(e)], as the charge density $|\psi({\bf r})|^{2}$ remains the same and concentrated around the I atoms. A similar behavior is observed for the conduction band minimum (CBM), where the DW terms remain similar while the positive FM term, leading to a CBM energy increase, is reduced by 71% due to local disorder.
In conclusion, we generalize the ASDM for the simultaneous treatment of anharmonicity and configurational entropy in both the inorganic and organic sublattices, as well as electron-phonon coupling, without long path-integral or classical MD runs. We clarify the origin of the gradual band gap blueshift of halide perovskites with temperature, attributing it to the low response of electrons to nuclear vibrations induced by positional polymorphism. We elucidate the influence of the A-site cation on local disorder, band gap, and electron-phonon renormalization, establishing rules for materials engineering. In the future, it will be interesting to investigate temperature-dependent properties of emerging hybrid materials with optoelectronic, ferroelectric, and thermoelectric applications Haque et al. (2020); Sidhik et al. (2022); Zhang et al. (2022); Dirin et al. (2023); Metcalf et al. (2023); Jiang et al. (2025). Our approach forms the basis for electron-phonon calculations in polymorphous solids, including phonon-assisted absorption Zacharias et al. (2015); Noffsinger et al. (2012); Zacharias and Giustino (2016), mobilities Herz (2017); Schlipf et al. (2018); Poncé et al. (2019); Cucco et al. (2024), polarons Franchini et al. (2021); René de Cotret et al. (2022); Lafuente-Bartolome et al. (2022); Zhang et al. (2023); Seiler et al. (2023); Biswas et al. (2024), and exciton-phonon interactions Alvertis et al. (2023); Dai et al. (2024b).
Data leading to the results of this study is available in the NOMAD repository Zacharias (2025).
Acknowledgements.
This research was funded by the European Union (project ULTRA-2DPK / HORIZON-MSCA-2022-PF-01 / Grant Agreement No. 101106654). Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Commission. Neither the European Union nor the granting authority can be held responsible for them. We thank D. R. Ceratti for graciously providing data on temperature-dependent band gaps of Br-based hybrid halide perovskite single crystals. J.E. acknowledges financial support from the Institut Universitaire de France. F.G. was supported by the Robert A. Welch Foundation under Award No. F-2139-20230405 and by the National Science Foundation under DMREF Grant No. 2119555. G.V. acknowledges funding from the ANR through the CPJ program and the SURFIN project (ANR-23-CE09-0001), the ALSATIAN project (ANR-23-CE50-0030), and the ANR under the France 2030 programme, MINOTAURE project (ANR-22-PETA-0015). We acknowledge that the results of this research have been achieved using computational resources from the EuroHPC Joint Undertaking and supercomputer LUMI [https://lumi-supercomputer.eu/], hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Extreme Scale Access call.